The College of Business at Northeast College is accumulating data as a first step in the preparation of next year's budget development. One cost that is being looked at closely is administrative costs as a function of student credit hours. Data on administrative costs and credit hours for the past thirteen months are shown below:

Month Administrative Costs Credit Hours
July $ 129,451 400
August 82,733 193
September 225,670 1,470
October 216,454 1,078
November 258,293 1,387
December 184,599 1,190
January 219,257 1,417
February 245,090 1,451
March 209,522 1,142
April 191,955 1,201
May 250,128 1,438
June 170,538 498
July 128,257 393
Total $ 2,511,947 13,258
Average $ 193,227 1,020


The controller's office has analyzed the data and has given you the results from the regression analysis:

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.9327181
R Square 0.869963033
Adjusted R Square 0.858141490
Standard Error 19,984.60564
Observations 13
ANOVA
df SS MS F Significance F
Regression 1 29,391,233,739 29,391,233,739 -73.5913299 3.34292E-06
Residual 11 4,393,229,090 399,384,462.7
Total 12 33,784,462,829
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 86,208.30382 13,651.03481 6.31514790 5.71542E-05 56,162.57879 116,254.0288 56,162.57879 116,254.0288
X Variable 1 104.9358161 12.23236465 8.578538912 3.34292E-06 78.01256307 131.859069 78.01256307 131.859069

  
Based on the results of the high-low analysis, the estimate of administrative costs in a month with 1,060 credit hours would be: (rounded to the nearest whole dollar)
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (225,670-82,733)/(1470-193)

= 111.93

Fixed cost = Total cost - Variable cost

= 225,670 - (1470*111.93)

= 61,133

Total Cost for 1060 hours

= (1060*111.93) + 61,133

= 179,779 (select the nearest answer)